a In how many different ways can it select a group of 4 members?




b In how many different ways can it elect a president, vice-president, secretary and treasurer? One member CANNOT hold more than one job. (optional)
1
Expert's answer
2020-08-06T17:15:05-0400

Solution:

a) Here n is not mentioned.so let us take n= 10 for example


You have a total of 10 people to choose from. Of these 10 people you are supposed to create a group of 4. Since order is not important here we use combination.

Formula:

nCr= "\frac{n!}{(n-r)!r!}"

Hence 10C4= "\frac{10!}{(10-4)!4!}"

="\frac{10!}{(6)!4!}"

= 210

Hence different ways to select a group of 4 members = 210


b)different ways can it elect a president, vice-president, secretary and treasurer= 10*9*8*7= 5040 ways